Show that x2i +2xyj - 4xzk is solenoidal


1
Expert's answer
2021-05-07T09:49:25-0400

Answer :-


"\overrightarrow{V}= x^2i +2xyj - 4xzk"


To prove that V is solenoidal


the condition for this is "\boxed{∇⋅\overrightarrow{V} =0}"


"\implies div.(\overrightarrow{v}) = 0"


"=\frac{∂(\overrightarrow{V})}{∂x} + \frac{∂(\overrightarrow{V})}{∂y} + \frac{∂(\overrightarrow{V})}{∂z}"


"=\frac{∂(x^2)}{∂x} +\frac{∂(2xy)}{∂y} +\frac{∂(-4xz)}{∂z}"


"= 2x +2x - 4x"

= 0


"∇⋅\overrightarrow{V}" comes out to be zero so ,

"\implies" given vector is  solenoidal